The treasurer of a large corporation wants to invest $20 million in excess short-term cash in a particular money market investment. The prospectus quotes the instrument at a true yield of 4.65 percent; that is, the EAR for this investment is 4.65 percent. However, the treasurer wants to know the money market yield on this instrument to make it comparable to the T-bills and CDs she has already bought. If the term of the instrument is 90 days, what is the discount yield (in percent) on this investment?
Answer #1

Discount yield = (F-P)/F * (360/T)

where F is the face value = $100

P is the price =$100/(1+0.0465) = 95.5566

T is the time in days = 90

Discount yield= (100-95.56)/100 * (360/90) = 0.1777 = 17.77%

Discount yield = 17.77%